# GraCoRe: Benchmarking Graph Comprehension and Complex Reasoning in Large Language Models

### v1

@@ Abstract
Evaluating the graph comprehension and reasoning abilities of Large Language Models (LLMs) is challenging and often incomplete. Existing benchmarks focus primarily on pure graph understanding, lacking a comprehensive evaluation across all graph types and detailed capability definitions. This paper presents GraCoRe, a benchmark for systematically assessing LLMs’ graph comprehension and reasoning. GraCoRe uses a three-tier hierarchical taxonomy to categorize and test models on pure graph and heterogeneous graphs, subdividing capabilities into 10 distinct areas tested through 19 tasks. Our benchmark includes 11 datasets with 5,140 graphs of varying complexity. We evaluated three closed-source and seven open-source LLMs, conducting thorough analyses from both ability and task perspectives. Key findings reveal that semantic enrichment enhances reasoning performance, node ordering impacts task success, and the ability to process longer texts does not necessarily improve graph comprehension or reasoning. GraCoRe is open-sourced at https://github.com/ZIKEYUAN/GraCoRe.
@@
Zike Yuan1,2, Ming Liu1, Hui Wang2,*, Bing Qin1,*, 1Harbin Institute of Technology, Shenzhen, China, 2Peng Cheng Laboratory, Shenzhen, China
@@ 1 Introduction
Graph understanding and complex reasoning are critical capabilities of Large Language Models (LLMs) due to their wide-ranging applications in fields such as social network analysis, drug discovery, recommendation systems, and spatiotemporal predictionsBrown et al. (2020). Thus, the ability to understand and reason about graph-structured data is particularly essential for advancing Artificial General Intelligence (AGI)Zhao et al. (2023). Graph-structured data primarily includes homogeneous and heterogeneous graphs. Research on homogeneous graphs often focuses on graph-structure-specific issues, such as protein-protein interaction predictionRao et al. (2014).In contrast to homogeneous graphs, pure graphs are simpler and do not contain attribute information for nodes and edge, such as graph-theoretic problemsBorgatti and Everett (2006). For heterogeneous graphs, tasks often rely on their rich semantic information, such as those related to knowledge graph reasoning. However, the use of LLMs for parsing, understanding, and reasoning about graph-structured data remains a challenging area of study.
@@
Firstly, the reasoning and comprehension capabilities of current large language models significantly deteriorate when handling complex graph-structured data with numerous nodes and edges. This degradation is partly due to the models’ difficulty in processing long textual inputs, especially when graph-structured data is described through lengthy texts. Long texts not only increase the computational burden but also introduce noise and redundant information, further diminishing the models’ ability to capture critical details. Secondly, textual descriptions of graph-structured data often involve complex entity relationships and abstract concepts, requiring models to both understand explicit information and infer implicit connections and relationships. Current research tends to rely on direct mappings from graph structure to answersWang et al. (2024), overlooking the potential to enhance the models’ deeper reasoning capabilities. Given the lack of systematic definitions and evaluation standards for the capabilities of large language models in graph understanding and reasoning, it is imperative to design a benchmark that extensively tests these abilities based on the models’ capabilities.
@@
In previous research, several benchmarks have been proposed to evaluate the understanding and reasoning capabilities of LLMs on graphs. However, existing benchmarks for evaluating the graph understanding capabilities of LLMs exhibit several limitations. Firstly, there is the issue of limited generalization: current benchmarks predominantly test either pure graphs or heterogeneous graphs in isolation, failing to provide a unified and systematic evaluation across both graph structures. Additionally, there is a lack of clear definition regarding model capabilities: traditional benchmarks for LLMs on graphs are primarily task-driven, inadequately assessing the specific abilities of LLMs on graph data. Hence, there is a need to develop better benchmarks that evaluate LLMs’ capabilities more comprehensively and from an ability-based perspective, specifically focusing on graph-structured data understanding and reasoning. Lastly, there is insufficient variety in model types and task categories: current benchmarks neither offer a clear classification of task types nor test a wide range of models.
@@
To address these challenges, we propose the GraCoRe benchmark, as shown in Figure 1, which aims to explore and evaluate the understanding and reasoning capabilities of mainstream LLMs. We have designed a three-tier hierarchical ability taxonomy that includes both capability-based and dataset-based categories. This taxonomy meticulously defines the model’s capabilities and ensures greater generalization in the testing range. Regarding model types and task divisions, our benchmark tests multiple existing closed-source and open-source models, and divides tasks into multiple dimensions based on model capabilities.
@@
Figure 2 illustrates the overall framework of our ability taxonomy. The first layer outlines two progressive overall capabilities: graph understanding and graph reasoning. Graph understanding is the most fundamental ability, reflecting the model’s proficiency in comprehending the nodes, edges, and overall structure of the graph within the contextual data. Built upon this foundation, graph reasoning encompasses the model’s ability to utilize graph-structured data to infer implicit information. The second layer categorizes the capabilities of LLMs into four types based on different data types. In the third layer, these four categories are further divided into 10 distinct capabilities, which are ultimately tested through 19 tasks. This taxonomy provides evaluation results across three levels, from general to detailed, allowing for the identification of deficiencies in models at varying levels of granularity. For each task, we meticulously design specific prompts and construct them based on predefined rules, ultimately textualizing the structural information. GraCoRe comprises a total of 11 datasets, consisting of 5,140 graphs. We control the complexity of these graphs by adjusting factors such as graph size and network sparsity.
@@
Finally, we conducted extensive experiments on GraCoRe to evaluate the graph understanding and reasoning abilities of large language models. We tested three closed-source LLMs and seven open-source LLMs. Our findings include the following:
  • We identified graph reasoning as a critical weakness in current LLMs. Most models cannot balance both graph understanding and reasoning, with GPT-4o being the most comprehensive model in both aspects.
  • LLMs perform better on graph reasoning tasks enriched with semantic information compared to tasks involving purely structural graph reasoning, indicating that textual information can enhance the graph reasoning abilities of LLMs.
  • Models exhibit high sensitivity to the ordering of nodes in textual graph data. An ordered naming of nodes can significantly improve model performance on graph tasks.
  • The capability of models to handle longer text inputs does not affect their performance in graph understanding and reasoning tasks, regardless of whether the graphs are complex with long textual descriptions or simple with short descriptions.
@@ 2 Related Work
LLMs for Graph Recent advancements in LLMs for graph tasks includes several notable contributions. Li et al. (2023) categorizes these tasks into three types: Enhancer, Predictor, and Alignment. Pan et al. provides a forward-looking roadmap for the unification of LLMs and Knowledge Graphs (KGs).Chai et al. (2023) proposes an end-to-end method for solving graph-related problems. Furthermore, Das et al. (2023) investigates methods to improve the zero-shot reasoning ability of LLMs over structured data in a unified manner. Yao et al. (2024) explores the graph generation capabilities of LLMs through systematic task designs and extensive experiments.
Benchmarks for LLMs on Graph Most benchmarks for evaluating LLMs on graph tasks are based on task testing. NLGraphWang et al. (2024) introduced a simple test dataset for eight graph tasks, while GPT4GraphGuo et al. (2023) tested LLM capabilities on semantic tasks. Liu and Wu (2023) assessed the capabilities of four LLMs in graph data analysis. Fatemi et al. (2023) proposed a method for describing graph data in text form. Perozzi et al. (2024) designed a hint method specifically for graph tasks. GraphInstructLuo et al. (2024) provided diverse processes and steps for graph data generation. HiGPTTang et al. (2024) proposed an evaluation method for heterogeneous graphs, and VisionGraphLi et al. (2024) assessed the capabilities of LLMs on image graphs.
@@ 3 GraCoRe
This section begins by describing a three-tier hierarchical taxonomy for large language models (LLMs) on graph data. Next, we explain the methodology used to collect the dataset. Finally, we present an analysis of the dataset’s statistics.
@@ 3.1 Hierarchical Ability Taxonomy
After analyzing Bai et al. (2024) evaluation of the LLMs in the multi-round dialogue task, we have developed a hierarchical taxonomy for classifying LLMs capabilities on graphs, essential for their evaluation. The taxonomy is structured into three levels, encompassing 19 tasks within 10 sub-capabilities. Table 1 summarizes each task with a brief description. This section elaborates on the three levels and their corresponding tasks, with detailed examples provided in the Appendix A.
@@
PureGra Bipartite Graph 460 18.52 48.9 1 1 5
Tree Structure Graph 460 19 18.49 1 1 5
Graph Traversal Graph 460 19 52.61 1 1 5
Shortest Path Graph 460 19 52.7 1 1 5
Max Flow Graph 460 19 105.72 1 1 1
Graph Coloring 460 19 52.7 1 1 1
Hamiltonian Graph 460 19 71.7 1 1 1
@@
Eulerian Graph 460 19.48 104.43 1 1 1
HeterGra IMDBText 500 27.21 30.32 3 4 4
ACMText 500 157.64 170.92 4 8 2
@@
Table 2: Benchmark Graph Statistics
@@ 3.1.1 Graph understanding
Understanding graph structure necessitates large language models capable of accurately answering questions about the graph’s basic properties and reconstructing its structural information from extensive text descriptions. This involves two core capabilities:
Pure graph understanding: Pure graphs refer to graph data containing only structural information, representing the simplest form of graph dataJin et al. (2023). The structural information of such graphs can be encapsulated in an adjacency matrix. Consequently, research on pure graphs often emphasizes the structural information. Assessing the ability of large language models to understand pure graphs should thus focus on their capacity to comprehend structural information. To this end, I have identified four sub-capabilities:
@@
  • Pure Graph Memory: This requires the large language model to reconstruct the input graph structure data, testing its memory capacity. This is specifically evaluated by the similarity scoreLahitani et al. (2016) of the reconstructed matrix.
@@
  • Graph Traversal: Traversal is fundamental to solving many graph theory reasoning problems. The model’s performance in reasoning tasks is influenced by its traversal capability. This study primarily tests whether the model can traverse a graph using Breadth-First Search (BFS)Beamer et al. (2013).
@@ 3.1.2 Graph reasoning
Graph semantic reasoning: Unlike structure-based reasoning, semantic information reasoning in graphs requires large language models to deeply understand the semantic meanings of nodes and edges, and to reason based on this semantic information. This involves modeling and reasoning about entities, relationships, and their interactions within the graph. Based on these tasks, we subdivide the semantic reasoning capabilities of large language models into Node Entity Reasoning and Link Relationship Reasoning. Corresponding tasks include node classification and link prediction. Previous studies have primarily addressed these two problems using graph neural networks (GNNs) such as GCNKipf and Welling (2016) and GraphSAGEHamilton et al. (2017). These methods typically require large structured graph data for training and cannot directly utilize graph data containing text information for inference. Consequently, investigating the use of large models for semantic reasoning is highly significant. This research will explore whether text enhancement impacts the performance of LLMs on these two tasks
@@ 3.2 Data Collection
We first divide the dataset into pure graphs and heterogeneous graphs to test the capabilities of large language models on these two types of datasets. For pure graphs, we customize unique data generation prompts based on the specific characteristics of each task, generating corresponding graph structure data using manually set rules. The scale of the graph is defined by the number of nodes and the sparsity of the network, ensuring that the generated data meets the specific needs of each task. For heterogeneous graph data, we use the ACMWang et al. (2019) and IMDBFu et al. (2020) datasets, converting them into text-based graph data. These are constructed according to a manually specified graph structure description framework, and unique prompts are designed for each task to build the dataset.
@@
GPT-4o 75.012 94.339 51.752 73.886 98.639 52.789 73.136 80.298 68.642 77.347 83.649
GPT-4 74.898 82.494 70.146 73.82 75.85 32.023 72.455 89.205 68.39 80.095 84.717
GPT-3.5 75.012 62.358 46.51 70.897 23.7 79.558 79.869 57.03 34.508 57.926 59.261
Llama3-ins-8b 69.065 32.597 32.881 56.348 69.277 30.239 79.718 39.398 67.886 73.499 44.663
Llama2-7b-chat 22.521 50.661 66.429 48.442 41.668 30.239 50.593 38.489 2.263 44.552 37.542
Chatglm3-6b 30.527 45.627 69.002 47.246 38.6 30.239 36.447 16.131 27.832 19.024 20.275
Chatglm2-32k-7b 18.69 27.563 22.492 7.851 48.241 100 16.249 64.211 27.832 19.024 31.312
Vicuna-v1.5-16k 40.933 38.372 0 47.446 29.397 52.14 30.698 46.851 66.123 68.003 38.343
Qwen2-7b-ins 68.322 14.238 71.289 65.582 50.432 62.037 37.885 53.849 67.886 23.849 76.707
Vicuna-v1.5-7b 24.523 51.253 69.002 7.984 23.7 30.239 22.452 14.041 68.138 36.185 23.034
Table 3: Standardized performance of graph understanding
@@
GPT-4o 99.268 41.714 81.963 61.792 80.692 58.327 86.743 79.702 1.01 1419.69
GPT-4 68.526 26.463 82.776 65.548 63.373 49.208 85.286 73.2 0.793 1318.473
GPT-3.5 39.607 32.563 80.845 82.338 48.218 76.353 75.344 62.262 0.419 1144.161
Llama3-ins-8b 58.365 47.815 38.887 91.617 20.074 85.048 46.462 70.62 0.226 1054.461
Llama2-7b-chat 50.55 81.369 23.444 29.095 24.404 65.75 35.835 43.583 -0.347 787.429
Chatglm3-6b 11.992 81.369 55.853 28.653 59.043 64.689 41.663 28.62 -0.421 752.831
Chatglm2-32k-7b 19.287 81.369 35.331 28.653 48.218 23.335 25.293 26.556 -0.595 671.508
Vicuna-v1.5-16k 43.776 20.362 24.257 28.653 20.074 8.49 15.694 42.551 -0.616 662.165
Qwen2-7b-ins 52.113 60.016 51.18 54.06 91.517 37.544 47.233 67.318 0.223 1053.057
Vicuna-v1.5-7b 56.021 26.463 24.968 29.095 43.889 30.758 39.949 5.091 -0.691 626.783
Table 4: Standardized performance of graph reasoning

### v2

@@ Abstract
Evaluating the graph comprehension and reasoning abilities of Large Language Models (LLMs) is challenging and often incomplete. Existing benchmarks focus primarily on pure graph understanding, lacking a comprehensive evaluation across all graph types and detailed capability definitions. This paper presents GraCoRe, a benchmark for systematically assessing LLMs’ graph comprehension and reasoning. GraCoRe uses a three-tier hierarchical taxonomy to categorize and test models on pure graph and heterogeneous graphs, subdividing capabilities into 10 distinct areas tested through 19 tasks. Our benchmark includes 11 datasets with 5,140 graphs of varying complexity. We evaluate four closed-source and eight open-source LLMs, conducting thorough analyses from both ability and task perspectives. Key findings reveal that OpenAI o1 model has amazing comprehension and reasoning capabilities, semantic enrichment enhances reasoning performance, node ordering impacts task success, and the ability to process longer texts does not necessarily improve graph comprehension or reasoning.GraCoRe is open-sourced at https://github.com/ZIKEYUAN/GraCoRe.
@@
Zike Yuan1,2, Ming Liu1,2, Hui Wang2,*, Bing Qin1,2,* 1Harbin Institute of Technology, Shenzhen, China, 2Peng Cheng Laboratory, Shenzhen, China {yuanzk,wangh06}@pcl.ac.cn {mliu,qinb}@ir.hit.edu.cn
@@ 1 Introduction
Graph understanding and complex reasoning are crucial capabilities of Large Language Models (LLMs), supporting applications in areas like social network analysis, drug discovery, recommendation systems, and spatiotemporal prediction (Brown et al., 2020). These abilities are particularly important for advancing Artificial General Intelligence (AGI) (Zhao et al., 2023). Graph-structured data mainly consists of homogeneous and heterogeneous graphs. Research on homogeneous graphs often addresses specific structural issues, such as protein-protein interaction prediction (Rao et al., 2014). Unlike homogeneous graphs, pure graphs are simpler, lacking node and edge attributes, as seen in graph-theoretic problems (Borgatti and Everett, 2006). For heterogeneous graphs, tasks leverage rich semantic information, like knowledge graph reasoning. However, enabling LLMs to effectively parse, understand, and reason with graph-structured data remains a significant challenge.
@@
Firstly, the reasoning and comprehension abilities of current LLMs degrade significantly when processing complex graph-structured data with numerous nodes and edges. This decline is partly due to challenges in handling lengthy textual inputs that describe such data. Long texts increase computational burden and introduce noise, reducing the models’ ability to capture essential details. Secondly, textual descriptions often involve complex entity relationships and abstract concepts, requiring models to understand explicit information and infer implicit connections. Current research primarily focuses on direct mappings from graph structure to answers (Wang et al., 2024), overlooking deeper reasoning capabilities. The absence of clear definitions and evaluation standards for graph reasoning highlights the need for a comprehensive benchmark to evaluate these abilities.
@@
Previous research has introduced several benchmarks to evaluate LLMs’ understanding and reasoning on graphs, but these benchmarks have notable limitations. Firstly, there is a limited generalization issue: most benchmarks test either pure or heterogeneous graphs separately, lacking a unified evaluation across both types. Additionally, there is a lack of clear definition of model capabilities: existing benchmarks are task-driven and fail to assess LLMs’ specific abilities with graph data. Therefore, more comprehensive, ability-based benchmarks are needed to evaluate understanding and reasoning on graph-structured data. Lastly, there is insufficient diversity in model types and tasks: current benchmarks do not clearly classify tasks or test a wide range of models.
@@
To address these challenges, we propose the GraCoRe benchmark, as shown in Figure 1, which aims to explore and evaluate the understanding and reasoning capabilities of mainstream LLMs. We have designed a three-tier hierarchical ability taxonomy that includes both capability-based and dataset-based categories. This taxonomy meticulously defines the model’s capabilities and ensures greater generalization in the testing range. Regarding model types and task divisions, our benchmark tests multiple existing closed-source and open-source models, and divides tasks into multiple dimensions based on model capabilities. In Appendix A, we compare in detail the differences between the existing benchmark and GraCoRe, as well as our advantages.
@@
Figure 2 presents the framework of our ability taxonomy. The first layer highlights two core capabilities: graph understanding and graph reasoning. Graph understanding reflects the model’s ability to comprehend nodes, edges, and graph structure within the context, while graph reasoning builds on this, focusing on inferring implicit information from graph-structured data. The second layer categorizes LLM capabilities into four types based on data types. In the third layer, these are further divided into 10 distinct capabilities, evaluated through 19 tasks. This taxonomy provides multi-level evaluation, enabling detailed identification of model weaknesses. Each task is designed with specific prompts, structured by predefined rules to textualize graph information. GraCoRe includes 11 datasets with 5,140 graphs, with graph complexity controlled by factors such as size and network sparsity.
@@
We conducted extensive experiments on GraCoRe to evaluate the graph understanding and reasoning abilities of LLMs, including four closed-source and eight open-source models. Key findings include:
  • Graph reasoning is a major weakness in current LLMs, with most models unable to balance understanding and reasoning. OpenAI o1 model has amazing understanding and reasoning capabilities.
  • LLMs perform better on graph reasoning tasks with semantic information than on purely structural tasks, showing that textual context enhances reasoning.
  • Model performance is sensitive to node ordering in textual graph data, where ordered naming improves results.
  • The ability to handle longer text does not impact performance, regardless of graph complexity or description length.
@@ 2 Related Work
LLMs for Graph Recent advancements in LLMs for graph tasks includes several notable contributions. (Li et al., 2023) categorizes these tasks into three types: Enhancer, Predictor, and Alignment. (Pan et al., 2023) provides a forward-looking roadmap for the unification of LLMs and Knowledge Graphs (KGs).(Chai et al., 2023) proposes an end-to-end method for solving graph-related problems. Furthermore, (Das et al., 2023) investigates methods to improve the zero-shot reasoning ability of LLMs over structured data in a unified manner. (Yao et al., 2024) explores the graph generation capabilities of LLMs through systematic task designs and extensive experiments.
Benchmarks for LLMs on Graph Most benchmarks for evaluating LLMs on graph tasks are based on task testing. NLGraph(Wang et al., 2024) introduced a simple test dataset for eight graph tasks, while GPT4Graph(Guo et al., 2023) tested LLM capabilities on semantic tasks. (Liu and Wu, 2023) assessed the capabilities of four LLMs in graph data analysis. (Fatemi et al., 2023) proposed a method for describing graph data in text form. (Perozzi et al., 2024) designed a hint method specifically for graph tasks. GraphInstruct(Luo et al., 2024) provided diverse processes and steps for graph data generation. HiGPT(Tang et al., 2024a) proposed an evaluation method for heterogeneous graphs, and VisionGraph(Li et al., 2024) assessed the capabilities of LLMs on image graphs.
@@ 3 GraCoRe
This section begins by describing a three-tier hierarchical taxonomy for LLMs on graph data. Next, we explain the methodology used to collect the dataset. Finally, we present an analysis of the dataset’s statistics.
@@ 3.1 Hierarchical Ability Taxonomy
After analyzing (Bai et al., 2024) evaluation of the LLMs in the multi-round dialogue task, we have developed a hierarchical taxonomy for classifying LLMs capabilities on graphs, essential for their evaluation. The taxonomy is structured into three levels, encompassing 19 tasks within 10 sub-capabilities. Table 1 summarizes each task with a brief description. This section elaborates on the three levels and their corresponding tasks, with detailed examples provided in the Appendix B.
@@
PureGra Bipartite Graph 460 19 49 1 1 5
Tree Structure Graph 460 19 18 1 1 5
Graph Traversal Graph 460 19 53 1 1 5
Shortest Path Graph 460 19 53 1 1 5
Max Flow Graph 460 19 106 1 1 1
Graph Coloring 460 19 53 1 1 1
Hamiltonian Graph 460 19 72 1 1 1
@@
Eulerian Graph 460 19 104 1 1 1
HeterGra IMDBText 500 27 30 3 4 4
ACMText 500 158 171 4 8 2
@@
Table 2: Benchmark Graph Statistics.
@@ 3.1.1 Graph understanding
Understanding graph structure necessitates LLMs capable of accurately answering questions about the graph’s basic properties and reconstructing its structural information from extensive text descriptions. This involves two core capabilities:
Pure graph understanding: Pure graphs refer to graph data containing only structural information, representing the simplest form of graph data(Jin et al., 2023). The structural information of such graphs can be encapsulated in an adjacency matrix. Consequently, research on pure graphs often emphasizes the structural information. Assessing the ability of LLMs to understand pure graphs should thus focus on their capacity to comprehend structural information. To this end, I have identified four sub-capabilities:
@@
  • Pure Graph Memory: This requires the large language model to reconstruct the input graph structure data, testing its memory capacity. This is specifically evaluated by the similarity score(Lahitani et al., 2016) of the reconstructed matrix.
@@
  • Graph Traversal: Traversal is fundamental to solving many graph theory reasoning problems. The model’s performance in reasoning tasks is influenced by its traversal capability. This study primarily tests whether the model can traverse a graph using Breadth-First Search (BFS)(Beamer et al., 2013).
@@ 3.1.2 Graph reasoning
Graph semantic reasoning: Unlike structure-based reasoning, semantic information reasoning in graphs requires large language models to deeply understand the semantic meanings of nodes and edges, and to reason based on this semantic information. This involves modeling and reasoning about entities, relationships, and their interactions within the graph. Based on these tasks, we subdivide the semantic reasoning capabilities of large language models into Node Entity Reasoning and Link Relationship Reasoning. Corresponding tasks include node classification and link prediction. Previous studies have primarily addressed these two problems using graph neural networks (GNNs) such as GCN(Kipf and Welling, 2016) and GraphSAGE(Hamilton et al., 2017). These methods typically require large structured graph data for training and cannot directly utilize graph data containing text information for inference. Consequently, investigating the use of large models for semantic reasoning is highly significant. This research will explore whether text enhancement impacts the performance of LLMs on these two tasks
@@ 3.2 Data Collection
We first divide the dataset into pure graphs and heterogeneous graphs to test the capabilities of large language models on these two types of datasets. For pure graphs, we customize unique data generation prompts based on the specific characteristics of each task, generating corresponding graph structure data using manually set rules. The scale of the graph is defined by the number of nodes and the sparsity of the network, ensuring that the generated data meets the specific needs of each task. For heterogeneous graph data, we use the ACM(Wang et al., 2019) and IMDB(Fu et al., 2020) datasets, converting them into text-based graph data. These are constructed according to a manually specified graph structure description framework, and unique prompts are designed for each task to build the dataset.
@@
OpenAI o1 57.836 81.591 58.491 58.959 96.545 71.257 66.263 72.029 58.731 62.463 75.531
GPT-4o 58.978 59.195 41.355 58.906 50.275 41.450 55.864 57.564 53.095 60.99 60.042
GPT-4 58.887 52.842 56.054 58.853 43.737 26.832 55.349 63.648 52.898 63.128 60.763
GPT-3.5 58.978 42.041 37.166 56.521 28.775 60.293 60.949 41.671 26.302 45.879 43.594
Llama3.1-ins-8b 45.869 57.845 35.643 36.175 33.679 27.517 43.522 52.908 50.327 34.094 44.614
Llama3-ins-8b 54.228 26.077 26.275 44.918 41.851 25.576 60.835 29.628 52.502 57.996 33.748
Qwen2-7b-ins 53.634 16.229 56.968 52.283 36.445 47.959 29.237 39.498 52.502 19.364 55.360
Llama2-7b-chat 17.048 35.766 53.084 38.613 33.930 25.576 38.836 29.007 0.991 35.473 28.946
Vicuna-v1.5-16k 31.755 29.175 0.000 37.818 30.410 40.993 23.808 34.718 51.118 53.720 29.486
Chatglm3-6b 23.442 33.066 55.140 37.659 33.050 25.576 28.151 13.735 21.061 15.610 17.300
Chatglm2-32k-7b 13.987 23.377 17.974 6.239 35.816 74.683 12.895 46.576 21.061 15.610 24.744
Vicuna-v1.5-7b 18.646 36.084 55.140 6.345 28.775 25.576 17.580 12.307 52.700 28.962 19.161
Table 3: Standardized performance of graph understanding.
@@
OpenAI o1 95.320 99.390 71.313 85.057 100.000 73.379 66.409 53.236 1.762 1403.799
GPT-4o 52.665 35.019 61.671 41.405 37.647 37.860 66.538 62.877 0.601 993.397
GPT-4 42.456 33.045 62.255 43.303 36.700 33.954 65.435 57.641 0.528 967.778
GPT-3.5 32.852 33.834 60.867 51.788 35.872 45.581 57.908 48.830 0.251 869.702
Llama3.1-ins-8b 35.101 33.045 35.446 54.021 33.624 77.194 40.064 50.908 0.115 821.597
Llama3-ins-8b 39.081 35.809 30.698 56.477 34.334 49.306 36.041 55.563 0.028 790.943
Qwen2-7b-ins 37.005 37.389 39.537 37.498 38.238 28.957 36.625 52.903 -0.038 767.630
Llama2-7b-chat 36.486 40.153 19.594 24.882 34.570 41.039 27.994 33.786 -0.524 595.774
Vicuna-v1.5-16k 34.236 32.255 20.179 24.659 34.334 16.512 12.746 32.955 -0.595 570.875
Chatglm3-6b 23.681 40.153 42.897 24.659 36.463 40.585 32.407 21.734 -0.608 566.369
Chatglm2-32k-7b 26.103 40.153 28.141 24.659 35.872 22.871 20.013 20.072 -0.765 510.845
Vicuna-v1.5-7b 38.303 33.045 20.690 24.882 35.635 26.050 31.109 2.783 -0.756 513.774
Table 4: Standardized performance of graph reasoning.